GlueX PID Meeting, May 28, 2010

From GlueXWiki
Jump to: navigation, search

Location, Time

Friday, May 28, 2010
2:00 pm EDT
JLab: CEBAF Center, Room F326/7
ESNet: 8542553
EVO: direct meeting link
Conference Group: 800-377-8846, participant code: 77438230
Slides: talks can be deposited in the directory /group/halld/www/halldweb/html/talks/2010-05 on the JLab CUE (you have to be a member of the "halld" group). This directory is accessible from the web at .


  1. Announcements
  2. Minutes from the last meeting
  3. Time-of-Flight
    1. Calibration: variation of pulser amplitude for time walk correction measurement
    2. Change in thickness of TOF counters?
    3. Status of PMT purchase (for prototype testing)
    4. Electronics: signal split scheme (no dynode signal)
    5. Prototype progress
    6. Status of the MOU
  4. Start Counter
    1. prototype progress: attenuation length measurements
  5. Target
    1. Status of the MOU
    2. Status of refrigerator plan
  6. Action Items


  • FSU: Paul Eugenio, Sasha Ostrovidov
  • JLab: Eugene Chudakov, Mark Ito (chair), Sascha Somov, Tim Whitlatch, Elliott Wolin, Beni Zihlmann
  • UMass: Rory Miskimen


  • Eugene announced that the recent change request approved has been approved and we can go ahead and buy some PMT's for the TOF prototype.
  • This generated some discussion:
    • Paul reported that Hamamatsu has offered to put a dynode output on the base for same price as for the standard configuration. He reminded us that the original plan was to use an XP2020 with both anode and dynode outputs.
    • Paul discussed using a passive splitter with Fernando Barbosa at the last collaboration meeting. We should discuss all options further.
    • Paul will send Mark a new quote for use in writing up the PR.

Reviewing minutes from the last meeting

We went over them.

Tim mentioned a recent meeting where there was a lab-wide review of helium demand in the future. At the meeting he mentioned the idea of getting a pulse tube refrigerator for target and our needs for the solenoid. Others at the meeting raised the possibility of supporting various other types of targets, include a FROST target. We agreed that for now our planning should only include hydrogen, deuterium and helium 4; other possibilities would have to be raised and developed before impacting current plans. The change request to support a purchase has been approved.


We went over several items.


Paul reminded us that in Hall B, time-walk studies are done with a laser pulsing system. To explore doing similar studies ourselves, he has received LED's from Nichia, both UV and blue. A circuit has been constructed to drive the LED's. The UV LED will be used to excite the scintillator. He is working with a scope right now. He can get down to 8 ns width and 5 V pulse height with the pulser he has.

We had originally discussed using the system being developed for the calorimeter to pulse the TOF. There was some concern about whether that system could vary output amplitude precisely enough to calibrate the time-walk.

change of thickness of counters

Paul proposed that we increase the thickness of the TOF counters to better match the size of the photocathode of the PMT, from 1 in. to 1-1/3 in. He has talked to Matt, who thought that there would be little effect on calorimeter. The width would remain the same at 6 cm.

There was some discussion about how much might really be gained. We probably should try to understand the limits to light collection in either plan.

wrapping material

Paul discussed choices of wrapping material:

  • He is now using Tyvek, some aluminized mylar, and aluminum foil
  • The material formerly known as VM2000 is now called ESR DF2000MA. It was originally sold as roll. Yuri Sharabian got some, PrimEx used it, the new start counter for CLAS used it, and they are planning on using it for the CLAS12 TOF.
  • It is not sold in rolls anymore; LCD screens is the current big market.
  • Paul has obtained a sample, 20 ft. long, 1.5 ft wide; he will fool around with it. He will also test other materials.
  • Cost: $1000 for a 150 ft. roll. The thickness is 2.6 mils.
  • It is non-conductive, non-magnetic, reflectivity north of 99%.
  • The current plan is to use black Tedlar for wrapping. Although DuPont says there is a shortage, they will send Paul a sample, 6 ft. by 150 ft., which ought to hold him for a while.

other items

  • prototype work: The scintillator has been machined down to the nominal 6 cm wide.
  • MOU: Paul has recent put effort on the MOU on hold pending completing of his DOE proposal. He will send Mark a final draft next week.


Rory has sent a recently revised version of the UMass MOU to JLab.

New Action Items

  1. Arrange a meeting to discuss signal splitting schemes for the TOF PMT signals.
  2. Think about light collection studies for the TOF.

Next Meeting

GlueX PID Meeting, June 11, 2010